Eight people were asked how many immediate family members they have. The results were: {5,2,4,4,3,6,7,7}. If one of the eight people was selected at random, what is the chance of selecting an even number or less than 4?
Thank you for your help!!

Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

even number= 4/8=1/2
under 4=2/8=1/4

Answer by solver91311(24713)



Of the choices given: 2, 4, 4, 3, and 6 are either even or less than 4, that is 5 choices. There are a total of 8 choices. So:

John